suppose $8000 is put into an account paying 6% compounded monthly. How long must the money be left in the account so that there is $20000 in it? Round your answer to the nearest tenth of a year
Answer by lwsshak3(11628)

suppose $8000 is put into an account paying 6% compounded monthly. How long must the money be left in the account so that there is $20000 in it? Round your answer to the nearest tenth of a year
***
compound interest formula: A=P(1+i)^n, P=initial investment, i=interest rate per compounding period, n=number of compounding periods, A=amount after n-periods
For given problem: P=8000, i=.06/12=.005, A=20000, compounding period=1 month
20000=8000(1+.005)^n
20/8=1.005^n
5/2=1.005^n
take log of both sides
log(5/2)=nlog(1.005)
n=log(5/2)/log(1.005)
n≈183.72 periods or≈15.3 yrs